Q.The equation x²/(10-a)+y²(4-a)=1 represents an ellipse if?

Grade:11

1 Answers

kkbisht
5 years ago
Dear student:
                    Equation of ellipse is x²/(10-a)+y²(4-a)=1 
                   Condition for a second degree equation in x & y given by  ax2 + 2hxy + by2 +2gx + 2fy +c=0 to represent a conic section (ellipse) if \Delta \neq0 and h2\neq af2 +bg2 +ch2 which is true but second condition h2
0\sqrt{}(10-a)/(4-a) or (10-a)/(4-a)>0 => case 1: 10-a>0 and (4-a) >0 => 10>a and 4>a =>  a
                                                             case 2; 10-a a>10 10
The values of   a10
kkbisht